# Ctrl-V: Higher Fidelity Video Generation with Bounding-Box Controlled Object Motion

### v2

@@ C.1 Stable Video Diffusion vs Box2Video: Generation Quality Comparison
C.2 BDD Result Visualizations: Bounding-Box Predictions and Motion-Controlled Video Generations
In the following section, we showcase a range of BDD generation results produced by our 3-to-1 generation pipeline in different scene scenarios, such as city, urban, highways, busy intersections and at night. Each visualization displays every 6th frame from a 25-frame clip, with the actual video generated at a frame rate of 5 fps. In the leftmost column labels, GT represents ground truth, PB represents predicted bounding-box frames, and PF represents predicted frames.
@@ C.2 BDD Result Visualizations: Bounding-Box Predictions and Motion-Controlled Video Generations
PF Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
PB Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
@@
PF Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
PB Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
@@
PF Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
PB Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Figure 12: 2D bounding-box frame predictions and motion-controlled video generations for various scenes.
@@
PF Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
PB Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
@@
PF Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
PB Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
@@
PF Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
PB Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
[…]
PF Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
PB Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Figure 13: 2D bounding-box frame predictions and motion-controlled video generations for various scenes.
C.3 vKITTI Result Visualizations: 3D Bounding-Box Predictions and Motion Controlled Video Generations
In the following section, we showcase a range of vKITTI generation results produced by our 3-to-1 generation pipeline in different scene scenarios. In Figure 14, we showcase our generations on the vKITTI test split. However, it is worth noting that the vKITTI test split comprises samples from only one type of scene (unseen during training). To provide a broader persepective, in Figure 15, we present additional results on the vKITTI train split. Each visualization displays every 6th frame from a 25-frame clip, with the actual video generated at a frame rate of 7 fps. In the leftmost column labels, GT represents ground truth, PB represents predicted bounding-box frames, and PF represents predicted frames.
@@
PF Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
PB Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
@@
PF Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
PB Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
@@
PF Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
PB Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Figure 14: 3D bounding-box predictions and motion-controlled video generations for various scenes on vKITTI test-split.
@@
PF Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
PB Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
@@
PF Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
PB Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
@@
PF Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
PB Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
[…]
PF Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
PB Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Figure 15: 3D bounding-box predictions and motion-controlled video generations for various scenes on vKITTI train-split.
C.4 KITTI Result Visualizations: Bounding-Box Predictions and Motion-Controlled Video Generations
In the following section, we showcase a range of KITTI generation results produced by our 3-to-1 generation pipeline in different scene scenarios, such as urban area, city streets and highway. Each visualization displays every 6th frame from a 25-frame clip, with the actual video generated at a frame rate of 7 fps. In the leftmost column labels, GT represents ground truth, PB represents predicted bounding-box frames, and PF represents predicted frames.
@@ C.4 KITTI Result Visualizations: Bounding-Box Predictions and Motion-Controlled Video Generations
PF Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
PB Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
@@
PF Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
PB Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
@@
PF Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
PB Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Figure 16: 3D bounding-box predictions and motion-controlled video generations for various scenes on KITTI test-split.
C.5 Visualizing Motion Controlled Generation via Box2Video
@@
Appendix D Key Insights and Future Directions
D.1 Noteworthy Discoveries in our Box2Video Model
@@
Our experiments have yielded findings suggesting that our Box2Video model not only understands bounding-box locations in scenes but also the encoded track IDs and 3D bounding-box orientation information in the rendered bounding-box frames. While these findings are not conclusive and require further investigation to be confirmed, they highlight potential directions for future projects. Here we share some of these findings to guide further research and development.
@@
    During preprocessing, we assign a random color to each track ID and fill the bounding boxes with their corresponding track IDs’ colors. When our Box2Video  receives unreasonable predictions, it still attempts to produce outputs that adhere to the given conditions, including the false bounding-box IDs.
@@
    In our 3D bounding box plots, we encode the vehicle’s orientation by marking the rear end with an “X”. Occasionally, our BBox Predictor incorrectly identifies the car’s orientation, marking the front side with an “X”. Therefore, during the video generation phase, we sometimes observe that the Box2Video model recognizes the wrong markings in the bounding box frames and generates a video of a car driving backwards.
@@
    We also observe that Box2Video may recognize the encoded object information regarding its class in the bounding box frames. Specifically, when an object is not present in the initial frame but its bounding box label appears in the final frame, the model can decode its class information and generate a correct instance of that class in the clip. However, we suspect that the object’s class information could also be inferred from the bounding box shapes. Further investigation is required.
@@
    Given the previous findings, we believe that plotting the bounding-box outlines a bit thicker to make them more prominent could potentially improve our results.
D.2 Future Directions
For future work, we believe it is worthwhile to develop a systematic approach to investigate the aforementioned observations. Additionally, it would be interesting to explore alternative methods for predicting bounding boxes. Furthermore, it would be beneficial to create a systematic method to measure the likelihood of the bounding box frames, instead of directly computing their IoU, recall, and precision rate with ground truth.

### v3

@@ E.1 Stable Video Diffusion vs Box2Video: Generation Quality Comparison
E.2 BDD Result Visualizations: bounding box Generations and Motion-Controlled Video Generations
In the following section, we showcase a range of BDD generation results produced by our 3-to-1 generation pipeline in different scene scenarios, such as city, urban, highways, busy intersections and at night. Each visualization displays every 6th frame from a 25-frame clip, with the actual video generated at a frame rate of 5 fps. In the leftmost column labels, GT represents ground truth, GB represents generated bounding box frames, and GF represents generated frames.
@@ E.2 BDD Result Visualizations: bounding box Generations and Motion-Controlled Video Generations
GF Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
GB Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
@@
GF Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
GB Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
@@
GF Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
GB Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Figure 14: 2D bounding box frame generations and motion-controlled video generations for various scenes.
E.3 vKITTI Result Visualizations: 3D bounding box Generations and Motion Controlled Video Generations
In the following section, we showcase a range of vKITTI generation results produced by our 3-to-1 generation pipeline in different scene scenarios. In Figure 15, we showcase our generations on the vKITTI test split. However, it is worth noting that the vKITTI test split comprises samples from only one type of scene (unseen during training). Each visualization displays every 6th frame from a 25-frame clip, with the actual video generated at a frame rate of 7 fps. In the leftmost column labels, GT represents ground truth, GB represents generated bounding box frames, and GF represents generated frames.
@@
GF Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
GB Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
@@
GF Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
GB Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
@@
GF Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
GB Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Figure 15: 3D bounding box generations and motion-controlled video generations for various scenes on vKITTI test-split.
E.4 KITTI Result Visualizations: bounding box Generations and Motion-Controlled Video Generations
In the following section, we showcase a range of KITTI generation results produced by our 3-to-1 generation pipeline in different scene scenarios, such as urban area, city streets and highway. Each visualization displays every 6th frame from a 25-frame clip, with the actual video generated at a frame rate of 7 fps. In the leftmost column labels, GT represents ground truth, GB represents generated bounding box frames, and GF represents generated frames.
@@ E.4 KITTI Result Visualizations: bounding box Generations and Motion-Controlled Video Generations
GF Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
GB Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
@@
GF Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
GB Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
@@
GF Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
GB Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Figure 16: 3D bounding box Generations and motion-controlled video generations for various scenes on KITTI test-split.
E.5 Nuscenes Result Visualization: bounding box Generations and Motion-Controlled Video Generation
@@
GF Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
GB Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
@@
GF Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
GB Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
@@
GF Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
GB Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Figure 17: 3D bounding box generations and motion-controlled video generations for various scenes on NuScenes.
E.6 Special Scenarios: Turning
In the following section, we showcase a range of generation results with special scenarios, such as turning (day/night). These results are produced by our 3-to-1 generation pipeline. Each visualization displays every 6th frame from a 25-frame clip, with the actual video generated at a frame rate of 7 fps. In the leftmost column labels, GB represents generated bounding box frames, and GF represents generated frames.
@@ E.6 Special Scenarios: Turning
GF Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
GB Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
@@
GF Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
GB Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
@@
GF Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
GB Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Figure 18: 3D bounding box generations and motion-controlled video generations for turning scenarios on NuScenes.
Frame 1 Frame 7 Frame 13 Frame 19 Frame 25
GF Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
GB Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
[…]
GB Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
[…]
GB Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
[…]
GB Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Figure 19: 2D bounding box generations and motion-controlled video generations for turning scenarios on BDD test-split.
E.7 Visualizing Motion Controlled Generation via Box2Video
@@
E.8 BDD100K Results under Varying Conditioning Scenarios
We investigate the impact of the number of conditioning bounding box frames on the quality and alignment of the output to the ground truth. We train distinct models for each conditioning configuration. Our findings are summarized in Table 7 and Table 8.
Key insights include:
  • Generation quality peaks when conditioned on the entire ground-truth bounding box trajectory.
  • Quality metrics show slight disagreement on the worst-performing setting, with conditioning on either: first+last bounding box frames or first three frames.
  • Comparing to the no-final-frame conditioning baseline, our results show that minimal conditioning with the central point of the final bounding box frame (termed ”trajectory frame”; see Figure 8) significantly enhances generation quality.
  • Table 8 demonstrates that incorporating the last bounding box frame significantly improves alignment between generated and ground-truth bounding box trajectories, as evidenced by the improved performance when conditioning on the first and last frames versus the first three frames.
  • Figure 23 showcases two predicted bounding box trajectories generated by BBox Generator without conditioning on last-frame bounding box locations. Notably, despite inaccurate last-frame bounding boxes compared to ground-truth, the BBox Generator still predict convincing bounding box trajectories independently of last-frame bounding box locations.
  • Figure 24 presents exemplary clips generated by the Box2Video model, conditioned on bounding box sequences derived without last-frame bounding box information.
  • Figure 23 and Figure 24’s visualization displays every 6th frame from a 25-frame clip, with the actual video generated at a frame rate of 7fps. In the leftmost column labels of Figure 24, GT represents ground-truth, GB represents generated bounding box frames, and GF represents generated frames.
  • Comparing the ground-truth and generated clips in Figure 24 reveals notable discrepancies. In the first example, a novel vehicle emerges (darker yellow bounding box), not present initially. This demonstrates BBox Generator’s ability to create new bounding box instances, which Box2Video then uses to generate corresponding visual content
[…]
BBox Generator + Box2Video 1-to-1 412.8 0.2967 0.5470 17.52
BBox Generator + Box2Video 3-to-1 373.1 0.3071 0.5407 17.37
BBox Generator + Box2Video 3-to-0 389.7 0.3085 0.5401 17.11
[…]
Teacher-forced Box2Video All 348.9 0.2926 0.5836 18.39
Table 7: Summary of generation quality metrics on BDD100K dataset: evaluating the impact of bounding box conditioning types on generation quality.
Method
# Cond.
BBox
[…]
(first+last)
[…]
(first+last)
[…]
(first+last)
BBox Generator 1-to-1 .$587\pm.214$$747\pm.187$$712\pm.194$$954\pm.047$$\textbf{.955}\pm.047$$9$002
3-to-1 .$\textbf{.647}\pm.176$$\textbf{.784}\pm.150$$\textbf{.783}\pm.156$$\textbf{.955}\pm.043$$\textbf{.955}\pm.042$$997\pm.001$
3-to-0 .$522\pm.204$$686\pm.204$$673\pm.201$$578\pm.214$$735\pm.222$$734\pm.201$
[…]
Table 8: Comparing trajectory alignment: comparing generated and ground-truth trajectory alignment under varying conditioning.
Frame 1 Frame 7 Frame 13 Frame 19 Frame 25
Figure 23: Bounding box trajectory generations are produced without conditioning on the last frame’s bounding box locations. The provided demos display generated trajectories with ground-truth bounding box trajectories overlaid for visual comparison.
Frame 1 Frame 7 Frame 13 Frame 19 Frame 25
GF Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
GB Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
[…]
GF Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
GB Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Figure 24: 2D bounding box frame generations and motion-controlled video generations on BDD100K test split where no final bounding box frame conditioning was provided.
Appendix F Trajeglish-Style Bounding Box Generator
To evaluate both the performance and simplicity of our bounding box generator model, we implemented a baseline method inspired by the Trajeglish model (Philion et al., 2023), which we refer to as the “Trajeglish-Style” model.
The original Trajeglish model employs a discrete sequence modeling approach to represent multi-agent road-user trajectories. Vehicle trajectories are modeled as a discrete sequence of actions, where each action corresponds to a token that describes the agent’s relative displacement. The model utilizes a GPT-like encoder-decoder architecture, taking as input the agents’ initial timestep information and map objects, and outputting a distribution over possible displacement actions. Trajeglish operates in bird’s-eye view (BEV) and achieved state-of-the-art (SOTA) performance on the 2023 WOMD Sim Agents Benchmark.
To adapt this method to our problem, we made several modifications, which are detailed below. Our “Trajeglish-Style” model focuses on predicting 2D bounding boxes from the ego perspective (POV). Unlike the BEV representation, these bounding boxes can overlap in 3D space and are subject to resizing during rollout. Additionally, the boxes can dynamically enter and exit the frame at any point in the sequence.
For our baseline, we define the set of possible actions as the displacement of a bounding box corner. Each corner’s displacement is represented by a 2D vector, where the vector’s magnitude is discretized uniformly into 16 values, ranging from 0.0 to 0.1 of the image size in the respective dimension. The direction of the displacement is discretized into 24 values, covering the full 360-degree range. This joint discretization yields a vocabulary size of 384 (16 magnitudes * 24 directions). At each timestep, the movement of a bounding box is determined by selecting two actions: one for the displacement of the top-left corner and one for the bottom-right corner.
As input to the model’s encoder, we provide the initial and final bounding box coordinates (1 or 3 frames), as well as the type of each agent (e.g., car, pedestrian, cyclist). In lieu of HDMap information (missing in most of the datasets used in this study) we use the initial natural image as contextual input to the encoder. The image is encoded using the same method as our diffusion-based bounding box generator (VAE image encoding combined with CLIP embeddings) to ensure fair comparison.
While the original Trajeglish model achieved SOTA performance in its intended task of predicting BEV agent trajectories, we do not claim that our “Trajeglish-Style” model reaches SOTA for the current task of ego POV 2D bounding box trajectory prediction. In fact, we consider this task to be significantly more challenging and less suited to the modeling approach of Trajeglish, which was not originally designed for the complexities predicting the motion of bounding boxes in the ego perspective, as outlined above. This increased difficulty accounts for the drop in performance, as seen in Table 1. Nevertheless, we believe that the “Trajeglish-Style” model serves as a valuable baseline for evaluating the performance of other models, including our own diffusion-based approach.
Despite the complexities inherent in this task, our diffusion-based BBox Generator, which operates directly on pixel images of bounding boxes, offers a compelling alternative due to its simplicity and strong performance.
The code for the “Trajeglish-Style” model implementation is available in the GitHub repository associated with this work.
Appendix G Key Insights and Future Directions
G.1 Failure Cases
Ctrl-V struggles to accurately encode and decode specific visual details, particularly road signs, building lettering, license-plate information and lane markings. This limitation stems from the constraints of the off-the-shelf VAE network used for image encoding and decoding. Figure 25 illustrates the VAE model’s difficulty in accurately encoding and decoding the ”MATTRESS FIRM” sign on the building. A potential direction for future research is exploring super-resolution techniques to recover fine-grained text details.
((a))
((b))
Figure 25: Failure case: VAE’s inability to reconstruct building signage.
Resolution degradation is a pervasive issue in video generation, primarily caused by error accumulation through temporal propagation. Notably, Ctrl-V exhibits this problem, particularly when the ego vehicle travels at higher speeds or changing directions. Figure 26 shows instances of degraded generation quality, where the last frame outputs suffer from a loss of realism. Future work can investigate multistage training paradigms, leveraging auxiliary networks to refine and enhance the visual quality of generated frames.
((a))
((b))
Figure 26: Failure case: resolution degradation over time.
G.2 Noteworthy Discoveries in our Box2Video Model
@@
Our experiments have yielded findings suggesting that our Box2Video model not only understands bounding box locations in scenes but also the encoded track IDs and 3D bounding box orientation information in the rendered bounding box frames. While these findings are not conclusive and require further investigation to be confirmed, they highlight potential directions for future projects. Here we share some of these findings to guide further research and development.
@@
    During preprocessing, we assign a random color to each track ID and fill the bounding boxes with their corresponding track IDs’ colors. When our Box2Video  receives unreasonable generations, it still attempts to produce outputs that adhere to the given conditions, including the false bounding box IDs.
@@
    In our 3D bounding box plots, we encode the vehicle’s orientation by marking the rear end with an “X”. Occasionally, our BBox Generator incorrectly identifies the car’s orientation, marking the front side with an “X”. Therefore, during the video generation phase, we sometimes observe that the Box2Video model recognizes the wrong markings in the bounding box frames and generates a video of a car driving backwards.
@@
    We also observe that Box2Video may recognize the encoded object information regarding its class in the bounding box frames. Specifically, when an object is not present in the initial frame but its bounding box label appears in the final frame, the model can decode its class information and generate a correct instance of that class in the clip. For example, observing Figure 27, we see the model successfully handle new objects entering the scene, accurately assigning the incoming car and pedestrian to their corresponding bounding boxes by leveraging encoded object IDs. However, we suspect that the object’s class information could also be inferred from the bounding box shapes. Further investigation is required.
@@
    Given the previous findings, we believe that plotting the bounding box outlines slightly thicker to make them more prominent could potentially improve our results.
Figure 27: A frame-by-frame visualization of Box2Video generation, conditioned on the ground-truth bbox frame sequence from the BDD dataset, demonstrating the model’s capability of handling incoming objects.
G.3 Future Directions
For future work, we believe it is worthwhile to develop a systematic approach to investigate the aforementioned observations. Additionally, it would be interesting to explore alternative methods for generating bounding boxes. Furthermore, it would be beneficial to create a systematic method to measure the likelihood of the bounding box frames, instead of directly computing their IoU, recall, and precision rate with ground truth.